OpenStudy (anonymous):

Rewrite the equation 4x - 2y = 8 in slope-intercept form. -2y= -4x+ 8 2y= 4x+ 8 y= 2x- 4 y= 2x+ 4

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Well tried but no..

OpenStudy (ghazi):

it's a

OpenStudy (anonymous):

slope intercept form is y=mx+b

OpenStudy (anonymous):

No.. not a ...

OpenStudy (ghazi):

-4x+8= -2y

OpenStudy (anonymous):

y=2x-4 here u can see http://www.wolframalpha.com/input/?i=4x+-+2y+%3D+8

OpenStudy (anonymous):

4x - 2y = 8 2y = 4x - 8 Or: y = 2x - 4..

OpenStudy (ghazi):

y=2x-4

OpenStudy (anonymous):

u can see the steps by looking at my link